OpenStudy (crashonce):

Does anybody know a validated way to prove 1+1=0? I mean not using 0/0 etc. WILL FAN AND MEDAL thanks! CO

OpenStudy (ikram002p):

u mean 1-1 =0 ?

OpenStudy (paki):

1+1=0 ikram

OpenStudy (ikram002p):

mmm in bolean function ?

OpenStudy (crashonce):

ya 1+1=0

OpenStudy (ikram002p):

well lets see in integer system ?

OpenStudy (crashonce):

idk i just wanna find a way of doing it as long as its legit

OpenStudy (ikram002p):

ok :) here is away to think of it , in Z group with adding operation to be a group it need to have 1_identity which is zero 2_and for any \(a \in Z \) there is "-a " such that a + -(a) = 0 ok so far ?

OpenStudy (ikram002p):

so we need to check if 1+1=0 from the above we know that 1+ (-1 )=0 right ?

OpenStudy (crashonce):

ya

OpenStudy (crashonce):

@ikram002p could you just type the method cos i need to put another question up thanks

OpenStudy (ikram002p):

thats it lol xD the Additive inverse is unique we have from hypothesis that 1-1 =0 and ur Qn is 1+1 =0 so 1 = -1 which is a contradiction :D

OpenStudy (crashonce):

ok thanks